OpenStudy (anonymous):

Sandra throws an object into the air with an initial vertical velocity of 38 feet per second, from a platform that is 30 feet above the ground. How long will it take the object to hit the ground?

OpenStudy (anonymous):

@hartnn

OpenStudy (cwrw238):

use the equations of motion for constant acceleration here we have a displacement (s) of -30 ft , initial velocity (u) of 38 ft/s and the acceleration (a) is that due to gravity which is -32 ft/s^2 and you need to find time. use s = ut + 0.5at^2 and solve for t

OpenStudy (anonymous):

alright ill give it a shot

OpenStudy (anonymous):

0.5at^2 + tu = -30

OpenStudy (anonymous):

nope im wronge ill try again

OpenStudy (anonymous):

Okay i'm missing something here. Could you show what you would do?

OpenStudy (cwrw238):

s = -30, u = 38 and a = -32 so we get -30 = 38t - 0.5*32*t^2 -30 = 38t - 16t^2 now solve this quadratic equation ffor t you'll have 2 solutions on positive and other negatrive the positive one will be the rquired time

OpenStudy (anonymous):

Okay, That clears it up Thank you for your help!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

It's 3!!

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

WhooHooo!

OpenStudy (anonymous):

you just got your self another fan!

OpenStudy (cwrw238):

there are 4 of these equations v = u + at s = ut + 0.5at^2 v^2 = u^2 + 2as ans s =2 (U +v) / t v = final velocity they are very useful for solving problems of motion under gravity